OpenStudy (gabylovesyou):

The table below shows the ticket rates for whale watching trips offered by Beluga Fun Tours. (wait for table)

OpenStudy (gabylovesyou):

Age Ticket Price Under three years free Three to twelve years $34 Over twelve years $45

OpenStudy (gabylovesyou):

On a certain day, the company took 103 people on whale watching trips. There were 64 children aged twelve and under, of which some children were under three years. If x represents the number of children under three years, which equation can be used to find the value of x, where C represents the total amount of money collected from tickets that day?

OpenStudy (gabylovesyou):

(64 − x)34 + (103 − 64) ⋅ 45 = C (103 + x)34 + (103 −(64 + x)) ⋅ 45 = C (103 − x)34 + (103 − 64) ⋅ 45 = C (64 + x)34 + (103 −(64 + x)) ⋅ 45 = C

OpenStudy (btaylor):

basically: (number of people over 12)($45) + (number of people ages 3-12)($34)=C so, we get: (103-64)($45) + (64-x)($34) = C so, the answer is A.

OpenStudy (gabylovesyou):

Thank you!!!

OpenStudy (btaylor):

you're welcome